### v1

# Construction of Functional Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

Abstract
The convergence of materials science and artificial intelligence has unlocked new opportunities for gathering, analyzing, and generating novel materials sourced from extensive scientific literature. Despite the potential benefits, persistent challenges such as manual annotation, precise extraction, and traceability issues remain. Large language models have emerged as promising solutions to address these obstacles. This paper introduces Functional Materials Knowledge Graph (FMKG), a multidisciplinary materials science knowledge graph. Through the utilization of advanced natural language processing techniques, extracting millions of entities to form triples from a corpus comprising all high-quality research papers published in the last decade. It organizes unstructured information into nine distinct labels, covering Name, Formula, Acronym, Structure/Phase, Properties, Descriptor, Synthesis, Characterization Method, Application, and Domain, seamlessly integrating papers’ Digital Object Identifiers. As the latest structured database for functional materials, FMKG acts as a powerful catalyst for expediting the development of functional materials and a fundation for building a more comprehensive material knowledge graph using full paper text. Furthermore, our research lays the groundwork for practical text-mining-based knowledge management systems, not only in intricate materials systems but also applicable to other specialized domains.
Introduction
In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions. These conventional research approaches significantly impede the life-cycle of high-performance material research. Given the specialization, inherent complexity, and vast knowledge base of material science, researchers focusing on a single direction often struggle to efficiently access and understand material knowledge from multidisciplinary studies. For instance, researchers in solar cell development might not fully comprehend studies related to solid-state batteries or organic light-emitting diodes. Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other. To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [1]. However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering. This fragmentation of knowledge represents a significant barrier to interdisciplinary collaboration and innovation. A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
Despite the existence of current databases of scientific literature such as Scopus, Web of Science, and Crossref, which offer ways to search for research papers based on specific labels, extracting useful information of material science from the vast ocean of literature remains demanding. To obtain clearer sense of materials properties some structured database project such as Materials Project [2], OQMD [3], and NOMAD [4] were developed. However, these databases contain many computational results obtained through techniques like Density Functional Theory (DFT) or Molecular Dynamics (MD) simulations [5]. While these computational databases can provide valuable references for predicting and understanding certain materials systems, they often face discrepancies with experimental observations. Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
Furthermore, the complexity of materials information extends beyond composition and structure to encompass their respective application fields. For instance, organic materials are commonly utilized in biological applications, semiconductors serve as integral components in electronics, and metals find applications in mechanical engineering. The process of designing novel materials typically commences with a clear understanding of their intended applications, aiming to maximize research efficiency. Given this context, compared to universal databases like the Materials Project, which might not be as useful due to their broad focus, specific databases focused on applications or traits hold the potential to provide more valuable information for researchers in relevant industries. This underscores the importance of developing dedicated databases that cater to the nuanced needs of materials science research, facilitating a more targeted approach to material discovery and application.
Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide. The adoption of knowledge graphs offers several advantages, including enhanced data interoperability, the ability to infer new knowledge through relational data analysis, and improved data quality and consistency through structured representation [6], [7]. These features make knowledge graphs particularly valuable for integrating diverse information sources and providing a unified view of a domain’s knowledge, thereby facilitating more informed decision-making and discovery [8]. However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [9]. This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [10]. With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [11]. For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [12]. Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [13]. A user-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [14]. Recently, a material knowledge graph, MatKG and MatKG2, containing information on material properties, structure, and applications, has been developed [1], [15].
However, these material knowledge graphs face even greater challenges. Firstly, although advancements in NLP technology have reduced the dependency on experts to a certain extent, training data still requires extensive annotation to enhance model accuracy [16]. Secondly, the construction of these knowledge graphs often involves predicting relationships between nodes to form triples, which means the entities represented in the KG are not always based on real instances [17]. This can diminish the authenticity and credibility of the KG. Additionally, this approach makes updating the knowledge graph difficult, as each new node introduced necessitates predicting its relationship with every other node, complicating the maintenance of a dynamic and accurate knowledge graph, especially in advanced fields like material science. Acknowledging these challenges, the emergence of LLMs like GPT and LLaMA represents a breakthrough, offering new solutions to enhance the extraction and credibility of structured information [18], [19]. The fine-tuning technique of LLMs can significantly enhance their performance in specific domain text tasks through training with fewer samples [20],[21]. This means improving the results of NER and RE without requiring a large amount of labor becomes possible and was adopt in our research.
In this paper, we have achieved significant advancements in the development of Functional Materials Knowledge Graph (FMKG), a pioneering graph database tailored for the field of functional materials. Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy. Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet. This method also makes updating KG very convenient. 2) We constructed the first knowledge graph dedicated for functional materials, where researchers can easily get the information about functional material through query the FMKG. 3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs.
Methods
Figure 1(a) illustrates the comprehensive workflow of our research. Through NER and RE tasks, we extract the structural information about catalyst, battery and solar cell. After ER and normalization, we integrate information from these three fields and construct a knowledge graph. Specifically, the pipeline displayed in the Figure 1(b), commencing with the manual annotation and normalization of the initial training data set, prompting the fine-tuned LLMs specifically for NER and RE tasks. This inference dataset is subsequently divided into ten batches, a crucial step for the iterative process that follows. Then, we finish ER task through the NLP technology including ChemDataExactor [22], mat2vec [23] and our expert dictionary. After ER, high-quality results are meticulously selected to augment the training set, thereby enhancing the model’s performance in subsequent iterations. Finally, the knowledge graph is constructed using the triples transferred from normalizd result after the last iteration.
Figure 1: (a) Workflow and (b) pipeline of the fine-tuned LLM for knowledge graph tasks.
Data preparation and schema design
Material experts undertook the annotation of nine distinct categories from 75 abstracts of research papers, act as LLM’s training dataset. These categories encompassed core labels - "Name," "Formula," or "Acronym" - which consistently represent specific materials across diverse papers, albeit in varied forms. Additionally, supplementary labels included "Descriptor," "Structure/Phase," "Application," "Property," "Synthesis," and "Characterization." The "Property" label served to delineate quantitative properties such as ’length,’ ’specific surface area,’ and ’mass,’ whereas the "Descriptor" label encapsulated qualitative attributes like ’stable,’ ’vertically,’ and ’safe.’ Since from the application aspects, developers always concern about its own properties, we also annotated the essential parameters as "Property" directly linked to specific "Application". For instance, attributes like ’specific capacity’ for lithium-ion batteries, ’energy conversion efficiency’ for solar cells, and ’H2 production rate’ for hydrogen evolution reaction were systematically annotated.
As the Figure 2 shows, the central node, "Material," is pivotal and connects to nodes that describe its nomenclature, composition, and various attributes. Specifically, "Material" is consisted of "Formula", "Name" and "Acronym". "Structure/Phase" describing its physical form; "Application" denoting its practical uses; and "Property" outlining its inherent characteristics. The "Material" node is also associated with "Descriptor" node that provide additional qualitative information. Moreover, the "Application" node branches out to "Property", "Descriptor" and a "Domain" suggesting further specification and contextual relevance of application. To save the source of nodes and relations, each node is connected to a "Digital Object Identifier (DOI)" node. If we want to determine the source article from which the relation is derived, we can obtain the result by querying the intersection of "DOI" node neighbors at both ends of the relation.
Figure 2: This schematic represents the FMKG schema (a) All the nodes in FMKG connect to all their sources - "DOI". (b) Relations between all labels. (c) Relations between core labels, the more arrows pointing to the label, the lower the priority it has.
Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction. This normalization ensured uniform representation of entities with similar meanings. For example, terms such as ’Lithium-Ion Battery’ and ’Li-ion batteries’ were standardized to ’lithium-ion battery,’ while phrases like ’solution casting method,’ ’solvent post-treatment method,’ and ’solution-based deposition’ were simplified to ’solution-processed.’ All annotated entities underwent this normalization process, ensuring consistency and facilitating effective training of LLMs. The field of functional materials encompasses a wide range of areas. At this stage, our priority is to focus on energy materials. We download 150,000 abstracts of peer-reviewed research articles of energy material science includes battery, solar cell and catalyst from the Web of Science. Each abstract was stored in a JSON file format, structured as "DOI - text," facilitating seamless processing and analysis.
LLMs training, evaluation and inference
The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA7b, LLaMA2 7b, and Darwin [24]. Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data. The models underwent training over 10 epochs with a batch size of 1. Additionally, 60 abstracts were annotated to assess the LLMs" performance. Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task"s ability to identify potential internal relations among relevant element sets. Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities. Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
Precision=True Positives (tp)True Positives (tp) + False Positives ($\text{Precision}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) +False Positives (fp)}}$ (1)
Recall=True Positives (tp)True Positives (tp) + False Negatives ($\text{Recall}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) + False Negatives (fn)}}$ (2)
$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
For NER, a true positive (tp) is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema. Conversely, a false positive (fp) occurs when the model incorrectly identifies an entity, relation, or standardization, and a false negative (fn) is when the model fails to identify a correct entity, relation, or schema element that should have been recognized. After definition, we can evaluate each task using Equations (1), (2), and (3).
After evaluation, we choose a fine-tuned LLM, which demonstrates optimal performance in both NER and RE task, to iteratively infer the 150,000 abstracts. The output of inferences are organized in the "DOI - text - response" format. Consequently, the fine-tuned LLM not only extracts entities but also assigns them to appropriate labels, thereby accomplishing NER and RE tasks concurrently. Moreover, every entity and relation identified in the response is traceable, enhancing the integrity and utility of the data.
Entity resolution
The quality of KG is crucial for its credibility, we need to further check and correct the inference results before graph construction. For the core label, we divide standardization into two steps: ER-NF/A to ER-N/F. In stage of ER-NF/A, we apply the ChemDataExtractor to derive the Formula and pairs of "Name - Acronym" from the abstract. Subsequently, we utilize the mat2vec model to perform word embedding on entities extracted by both the LLM and ChemDataExtractor. Through comparing their similarities, we aim to correct the core entities to guarantee their precision and the accurate correspondence between "Name" and "Acronym". However, The Name and Formula of materials are often placed in incorrect labels and are difficult to distinguish through ChemDataExtractor. Therefore, we conduct ER-N/F. We selected 2000 correctly classified labels as the training set and used to fine-tune LLM to complete a binary classification problem. Additionally, we selected other 200 labels to evaluate the performance of binary classification.
To normalize entities and relations in other label (ER-OL), we conduct a Word Embedding Clustering to create a expert dictionary. We convert words into vector form using mat2vec and group semantically similar words into the same cluster based on their vector similarity. Specifically, we develop a Density-Based Dynamic Vector Clustering method that dynamically forms clusters based on the similarity of vector representations, without predefining the number of clusters. This approach utilizes a similarity threshold to determine cluster membership, allowing entities to join existing clusters or form new ones based on their proximity in the feature space, effectively adapting to the inherent structure of the data.
By naming each cluster through experts in the field of material science, we extracted a total of approximately 600 vocabulary, including the structure, phase, applications, characterization methods, synthesis methods, properties, and descriptions of most energy materials. Subsequently, we adopt analytic strategies and embed the vocabulary from expert dictionary and the enetities extracted by LLM using mat2vec. After similarity comparison, we finish the standardization and ensure that both entities and relations are correct. It is worth mentioning that the content of "Property" and "Descriptor" is often not limited to the energy field and may cover more information. Therefore, we implement relatively tolerant standardization for these two labels, which only normalize them with matching content but does not delete unmatched entities and relations.
To improve the caliber of the training dataset, we select a small number of high-quality data from the normalized inference output of each iterative process, incorporating this data into the training set. High-quality data is characterized by its high precision and recall metrics. Furthermore, the performance of the fine-tuned LLM, trained on each progressively augmented training set, is rigorously evaluated to ensure optimal effectiveness.
Knowledge graph construction
To structure the inference results into triples, we employ three labels that signify the material as the core label, with the core label serving as the head, the names of other labels forming the relations, and their values acting as tails. Within the hierarchy of core labels, "Formula" is accorded the highest priority, followed by "Name", and then "Acronym". Furthermore, each head and tail node is linked to the DOI of the source article associated with the triplet. This setup allows us to ascertain the provenance of the relation between any two nodes by examining the intersection in their connected entities. Then we transfer the triples into FMKG and store the FMKG via graph database Neo4j, which also supports the subgraph matching function, where subgraph matching naturally suits the need to search for certain materials with user-input conditions. To facilitate access to the detailed information, we also make the dataset available in the CSV format for straightforward data handling.
Result
In Table 1, the performance comparison of LLMs includes Darwin, LLaMA, and LLaMA2 in NER, RE and ER task within the materials science domain is presented. Darwin achieves markedly higher F1 scores in both NER and RE tasks than the LLaMA 7b and LLaMA2 7b models. This indicates that Darwin can perform more satisfactory results in text tasks in the field of materials science. However, there has been no significant difference in their effectiveness for ER tasks between these models. We speculate that this is due to insufficient contextual memory capabilities of the LLMs.
In response to the challenge, entity resolution was applied to the inference outputs, which substantially improved the performance in Standardization (ER) tasks. This refinement process has also led to enhancements in the NER and Relation RE tasks as well. The improvements can be attributed to the ER task removing some of the incorrect inferences that resulted from hallucinations and providing a secondary adjustment of the extracted relations.
Model Task Precision Recall F1 score
NER 0.6101 0.6216 0.6158
Llama 7b RE 0.5305 0.5405 0.5355
ER 0.3687 0.3757 0.3722
NER 0.7419 0.7667 0.7541
Llama2 7b RE 0.6452 0.6667 0.6557
ER 0.4484 0.4633 0.4557
NER 0.8013 0.7935 0.7974
Darwin RE 0.7036 0.6968 0.7002
ER 0.4593 0.4548 0.4571
NER 0.9520 0.9083 0.9296
Darwin (Normalized) RE 0.9039 0.8625 0.8827
ER 0.9127 0.8708 0.8913
Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
To gain a more comprehensive understanding of the improvements brought by each step in the standardization process, we conducted ablation experiments, the results are shown in Table 2 which clearly shows that each method contributes to the improvement of ER. The result indicates the most effective method is to optimize through the expert dictionary, which indicates that in addition to the accuracy of the expert dictionary, it can also fully cover all other labels except for the three labels representing the material. The improvement of ER-NF/A is also significant, as this process can eliminate most erroneous materials, and even has a stronger improvement on NER than ER-OL.
  Method NER F1 $\nabla$ RE F1 $\nabla$ ER F1 $\nabla$
Darwin (Normalized) 92.96 - 88.27 - 89.13 -
ER-N/F 92.96 - 83.29 -4.98 89.13 -
ER-NF/A 84.01 -8.85 83.07 -5.20 84.54 -4.59
ER-OL 88.59 -4.37 80.20 -8.07 50.00 -38.83
Table 2: Result of the ablation experiment in normalization.
Besides, ER-N/F pay attention to adjust the relation error between "Name" and "Formula", from a contribution perspective, it is also an indispensable part of ER. We also evaluated the LLM classifier by selecting 200 manually labeled data, and the statistical evaluation results in Table 3 shows the strong performance of LLM used in classification task.
Model Accuracy F1 Score Specificity
Darwin 0.9650 0.9581 0.9741
Table 3: Binary classification performance through Fine-tuned Darwin.
The triples are transferred from normalized data to construct the FMKG which contains 162,605 nodes and 731,772 edges, the schematic diagram shows in Figure 3. In the graph, there are 11 types of nodes and 13 types of relations. The FMKG store large amount of information in functional materials includes battery, catalyst and solar cell. For example, as the Figure 4 shows, by analyzing the frequency of material appearances in the battery domain, we observe that Co2O3 is the most frequently occurring material, followed by MoS2, graphite, TiO2, LiCoO2 etc which are commonly used in the battery domain. Likewise, it is evident that the most prevalent application within the battery field is lithium-ion batteries, which appear with a frequency significantly higher than other applications in the battery domain. These insights align with our understanding of these domains and materials, confirming that our knowledge graph effectively stores factual knowledge and offers a knowledge platform for material researchers.
Figure 3: (a) Global schematic diagram of FMKG; (b) Local schematic diagram of FMKG.
Figure 4: The image shows the top ten materials and applications in each field.
For further validation of FMKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science. After splitting these triples apart, annotator can obtain 1000 entities and 500 relations. The report from annotator shows in Table 4.
“Name” is the highest priority core label and only serves as the head in the knowledge graph, so it will not be used as a relation. Notably, the labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation. This is due to the rigorous standardization of entities under these categories, using our expert dictionary. In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation. But the results are still satisfactory.
"Name" and "Acronym" has also achieved a commendable 100% entity accuracy. However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations. A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor. The impact of this kind of error is not significant, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
Label Number in total Entity disagree % disagree Relation disagree % disagree
Formula 257 0 0 N/A N/A
Name 227 0 0 12 5.3
Acronym 26 0 0 3 11.5
Descriptor 127 7 5.5 12 9.4
Property 147 12 8.2 9 6.1
Application 73 0 0 0 0
Structure/Phase 36 0 0 0 0
Synthesis 50 0 0 0 0
Characterization 57 0 0 0 0
Table 4: Human evaluation metric on randomly selected triples.
Discussion
In this research, we construct a functional material KG by fine-tuned LLMs and ensuring the traceability of all information. During the process, we encountered two main challenges and proposed corresponding solution: 1) Fine tuning LLM reduces the requirement for training set volume, but requires higher quality for each data in training set. when the quality of training dataset is limited, the performance of NER and RE is not good. Recursively generating training set data automatically from inference result is a good solution. High recall and precision data from each inference can be put into the training set after simple manual check. 2) Fine-tuned LLMs perform well on NER and RE task, but for ER task, they can not shows a same. However, the ER is also important in graph construction. To address the issue, we apply several ER methods and technologies including ChemDataExtractor, mat2Vec, Word Embedding Clustering and construct an expert dictionary storing information about energy material. Theses methods and technologies can not only achieve high perfromance ER task, but also remove some erroneous entities and relations.
To better demonstrate the reliability of our method and FMKG, we compare them with existing structural information extraction method and material KG. We employ a similar approach to NERRE task of John Dagdelen et al. [21] to extract structured information from the text. The difference is that we have applied ER method to correct entities and relations, and finish ER tasks before graph construction. We evaluated our pipeline on NERRE’s evaluation dataset. It worth to mention that cross dataset evaluation allow for flexibility of multiple plausible annotations of an entity under different labels (e.g., "cathode" in NERRE is placed in "Application" label while our expert tend to place it in "Descriptor" since "cathode" and "anode" are component of "battery" which is an application.), and presented the results in a Table 5. From the results, it is evident that our pipeline has slightly lower recall across labels compared to the NERRE task, but our method holds a competitive edge in precision. This is because we believe that the primary goal of KG construction is to ensure the accuracy of the data. Hence, in the process of ER, our strategy tends to sacrifice a portion of recall in exchange for higher precision. However, when considering both recall and precision, our method scores higher in F1 across the "Acronym", "Application", "Structure/Phase", and "Descriptor" labels compared to NERRE, with the improvement in "Acronym" being particularly significant due to its lower precision in NERRE. This underscores the advantage of our ER pipeline and indirectly attests to the credibility of the FMKG.
Label Extraction recall (NERRE) Extraction precision (NERRE) Extraction F1 (NERRE) Extraction recall (DarwinN) Extraction precision (DarwinN) Extraction F1 (DarwinN)
Formula 0.943 0.943 0.943* 0.828 1.000 0.906
Name 0.692 1.000 0.818 0.667 1.000 0.800
Acronym 0.667 0.400 0.500 0.571 1.000 0.727
Descriptor 0.576 0.905 0.704 0.781 0.962 0.862
Application 0.797 0.870 0.832 0.750 1.000 0.857
Structure/Phase 0.754 0.920 0.829 0.613 0.905 0.730
*Best F1 scores for each label.
Table 5: Evaluation comparison of NERRE and DarwinN.
Knowledge graphs specifically dedicated to general materials are rare. In our discussion, we draw a comparison with MatKG2 [15], showcasing the differences in construction methodologies between MatKG2 and FMKG, as depicted in the Figure 5. The construction of MatKG2 involves in a multi-step NER and RE pipeline, which can not catch the source of relations. In our method, we use end-to-end method that use a single LLM have been investigated for joint named entity recognition and relation extraction. It can not only catch keep source of each triple in KG so that enhance the factual basis of KG, but also perform better on RE task.
Figure 5: Schematic comparison of FMKG and MatKG2.
KGs are always dynamic, especially in cutting-edge field such as material science. The LLM internalizes a large amount of knowledge during the training process, it can perform good inference even when new nodes and relationships appear, making KG easy to update. Besides, our work is a starting point where other material researchers can continue build FMKG using our result as training dataset, using our dictionary, ChemDataExtractor and mat2vec to achieve ER task, our method can also be extend to full paper, extract more functional material information to construct a larger and more comprehensive functional material knowledge graph. Besides, our pipeline can also be used in other specialized domains to construct domain-specific KGs.
Conclusion
In this study, we introduce a new NLP pipeline for KG construction, which aim to efficiently extract the triples from unstructured scientific texts. The main feature of the method is that it can fine-tune the LLMs by annotating a small amount of data, and use the fine-tuned LLM to extract structured information from a large amount of unstructured text. The entire process does not rely on any prediction, which can maximize the authenticity and traceability of structured information. By employing this method, we construct a Functional Material Knowledge Graph (FMKG) contains the materials and their related knowledge from abstract of 150,000 peer-reviewed paper. After analyzing, we have demonstrated the effectiveness and credibility of FMKG.
In addition, our method and KG have great potential in different dimensions. Firstly, enhancing the depth of structured information extraction to encompass entire research papers promises a richer, more detailed knowledge graph. This involves not only expanding the scope of data analyzed but refining the process to capture nuances within complex scientific texts. Secondly, refining entity labels within our system allows for a more precise categorization of data, including incorporating detailed attributes such as synthesis conditions or property parameters, which significantly improves the granularity and utility of the knowledge graph. Thirdly, the versatility of our NLP pipeline suggests its applicability across different scientific domains, offering a template for constructing domain-specific knowledge graphs beyond material science. Lastly, integrating FMKG with existing knowledge graphs like MatKG opens avenues for creating a more interconnected and comprehensive dataset, facilitating advanced research and application development in material science and beyond.