### v2

# Construction and Application of Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

Abstract
Knowledge in materials science is widely dispersed across extensive scientific literature, posing significant challenges for efficient discovery and integration of new materials. Traditional methods, often reliant on costly and time-consuming experimental approaches, further complicate rapid innovation. Addressing these challenges, the integration of artificial intelligence with materials science has opened avenues for accelerating the discovery process, though it also demands precise annotation, data extraction, and traceability of information. To tackle these issues, this article introduces the Materials Knowledge Graph (MKG), which utilizes advanced natural language processing techniques, integrated with large language models to extract and systematically organize a decade’s worth of high-quality research into structured triples, contains 162,605 nodes and 731,772 edges. MKG categorizes information into comprehensive labels such as Name, Formula, and Application, structured around a meticulously designed ontology, thus enhancing data usability and integration. By implementing network-based algorithms, MKG not only facilitates efficient link prediction but also significantly reduces reliance on traditional experimental methods. This structured approach not only streamlines materials research but also lays the groundwork for more sophisticated science knowledge graphs.
Introduction
In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions. These conventional research approaches significantly impede the life-cycle of high-performance material research. Given the specialization, inherent complexity, and vast knowledge base of material science, researchers focusing on a single direction often struggle to efficiently access and understand material knowledge from multidisciplinary studies. For instance, researchers in solar cell development might not fully comprehend studies related to solid-state batteries or organic light-emitting diodes. Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other. To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [1]. However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering. This fragmentation of knowledge represents a significant barrier to interdisciplinary collaboration and innovation. A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
Despite the existence of current databases of scientific literature such as Scopus, Web of Science, and Crossref, which offer ways to search for research papers based on specific labels, extracting useful information of material science from the vast ocean of literature remains demanding. To obtain clearer sense of materials properties some structured database project such as Materials Project [2], OQMD [3], and NOMAD [4] were developed. However, these databases contain many computational results obtained through techniques like Density Functional Theory (DFT) or Molecular Dynamics (MD) simulations [5]. While these computational databases can provide valuable references for predicting and understanding certain materials systems, they often face discrepancies with experimental observations. Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide. The adoption of knowledge graphs offers several advantages, including enhanced data interoperability, the ability to infer new knowledge through relational data analysis, and improved data quality and consistency through structured representation [6], [7]. These features make knowledge graphs particularly valuable for integrating diverse information sources and providing a unified view of a domain’s knowledge, thereby facilitating more informed decision-making and discovery [8]. However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [9]. This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [10]. With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [11]. For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [12]. Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [13]. A user-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [14]. Recently, a material knowledge graph, MatKG and MatKG2, containing information on material properties, structure, and applications, has been developed [1], [15].
However, these material knowledge graphs face even greater challenges. Firstly, although advancements in NLP technology have reduced the dependency on experts to a certain extent, training data still requires extensive annotation to enhance model accuracy [16]. Secondly, the construction of these knowledge graphs often involves predicting relationships between nodes to form triples, which means the entities represented in the KG are not always based on real instances [17]. This can diminish the authenticity and credibility of the KG. Additionally, this approach makes updating the knowledge graph difficult, as each new node introduced necessitates predicting its relationship with every other node, complicating the maintenance of a dynamic and accurate knowledge graph, especially in advanced fields like material science. Acknowledging these challenges, the emergence of LLMs like GPT and LLaMA represents a breakthrough, offering new solutions to enhance the extraction and credibility of structured information [18], [19]. The fine-tuning technique of LLMs can significantly enhance their performance in specific domain text tasks through training with fewer samples [20],[21]. This means improving the results of NER and RE without requiring a large amount of labor becomes possible and was adopt in our research.
In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science. Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy. Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet. This method also makes updating KG very convenient. 2) We constructed the first accurate knowledge graph dedicated for materials, where researchers can easily get the information about material through query the MKG. 3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs. 4) We demonstrate a similarity calculation method based on Jaccard Similarity for materials and applications.
Methods
Figure 1(a) presents the elaborate workflow of our study. Initially, through Named Entity Recognition (NER) and Relation Extraction (RE) tasks, we extract structural information pertinent to catalysts, batteries, and solar cells. Subsequent Entity Resolution (ER) and data normalization steps enable the integration of information from these distinct fields into a unified knowledge graph. Specifically, as outlined in Figure 1(b), the process begins with the manual annotation and normalization of the initial training dataset, essential for fine-tuning Large Language Models (LLMs) tailored for NERRE tasks. The inference dataset is then methodically divided into ten batches, facilitating the iterative refinement process that follows. The ER tasks are conducted using advanced NLP technologies, including ChemDataExtractor [22], mat2vec [23], and our expertly curated dictionary. After ER, we selectively enhance the training dataset with high-quality results to improve the model’s performance in subsequent iterations. The knowledge graph is finally constructed using the normalized data from the last iteration. To complete the graph and predict potential material applications, we employ both network-based algorithms and graph embeddings. This methodology provides critical insights and recommendations for researchers in the materials science domain.
Figure 1: (a) Workflow and (b) pipeline of the fine-tuned LLM for knowledge graph tasks.
Data preparation and schema design
Material experts annotated nine distinct categories from the abstracts of 75 research papers, forming the training dataset for the Large Language Models (LLMs). The structure of the data reflects the structure of the KG. As illustrated in Figure 2, the "Material" node serves as the central hub within the graph, linking to nodes that encapsulate its nomenclature, composition, and various attributes. Specifically, the "Material" node comprises sub-nodes such as "Formula," "Name," and "Acronym" for identifying materials. The "Structure/Phase" node describes the material’s physical state, while the "Application" node denotes its practical uses, and the "Property" node details its inherent characteristics. Additionally, the "Descriptor" node provides qualitative information enhancing the contextual understanding of the material.
Furthermore, the "Application" node is extended to include "Property", "Descriptor", and "Domain" nodes, indicating the specific attributes and the broader context of the material’s application. To maintain data provenance and traceability, each node is linked to a "Digital Object Identifier (DOI)" node. This allows for source verification where querying the intersections of "DOI" node connections at both ends of a relation can pinpoint the source article from which the relations was derived, ensuring data integrity within the graph structure.
Figure 2: This schema represents the ontology of MKG (a) All the nodes in MKG connect to all their sources - "DOI". (b) Relations between all labels. (c) Relations between core labels, the more arrows pointing to the label, the lower the priority it has.
Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction. This normalization ensured uniform representation of entities with similar meanings. For example, terms such as "Lithium-Ion Battery" and "Li-ion batteries" were standardized to "lithium-ion battery", while phrases like "solution casting method", "solvent post-treatment method", and "solution-based deposition" were simplified to "solution-processed", ensuring consistency and facilitating effective training of LLMs. The field of materials encompasses a wide range of areas. At this stage, our priority is to focus on energy materials. We download 150,000 abstracts of peer-reviewed research articles of energy material science includes battery, solar cell and catalyst from the Web of Science. Each abstract was stored in a JSON file format, structured as "DOI - text," facilitating seamless processing and analysis.
LLMs training, evaluation and inference
The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA 7b, LLaMA2 7b [24], and Darwin [25]. Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data. The models underwent training over 10 epochs with a batch size of 1. Additionally, 60 abstracts were annotated to assess the LLMs" performance. Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task"s ability to identify potential internal relations among relevant element sets. Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities. Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
For NER, a true positive is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema. Conversely, a false positive occurs when the model incorrectly identifies an entity, relation, or standardization, and a false negative is when the model fails to identify a correct entity, relation, or schema element that should have been recognized. After defining these terms, we can evaluate each task using standard precision, recall, and F1 score metrics.
After evaluation, we choose a fine-tuned LLM, which demonstrates optimal performance in both NER and RE task, to iteratively infer the 150,000 abstracts. The output of inferences are organized in the "DOI - text - response" format. Consequently, the fine-tuned LLM not only extracts entities but also assigns them to appropriate labels, thereby accomplishing NER and RE tasks concurrently. Moreover, every entity and relation identified in the response is traceable, enhancing the integrity and utility of the data.
Entity resolution
The quality of KG is crucial for its credibility, to further check and correct the inference results before graph construction. To ensure the precision of these results, we initially employ ChemDataExtractor to identify chemical formulas and "Name-Acronym" pairs from abstracts. Subsequently, entities recognized by both the Large Language Model (LLM) and ChemDataExtractor are embedded using the mat2vec model. We analyze their similarities to rectify core entities and ensure accurate "Name" to "Acronym" associations. Given the frequent mislabeling of "Name" and "Formula", we progress to the Entity Resolution - Name/Formulas (ER-N/F) phase. Here, we refine the LLM using a specifically curated training set comprising 2,000 accurately labeled entities for binary classification, which is evaluated against an additional set of 200 labels.
For other labels, we implement a Density-Based Spatial Clustering of Applications with Noise [26] algorithm to create the expert dictionary (ER-ED). This algorithm dynamically forms clusters based on vector similarity, without the need to predefine the number of clusters. Each cluster is named by material science experts, leading to the creation of an expert dictionary containing approximately 600 terms related to structures, phases, applications, and more. This dictionary, along with the entities extracted by the LLM, undergoes similarity analysis to standardize and confirm the accuracy of both entities and relations. To elevate the quality of the training dataset, we selectively integrate high-quality data from normalized inference outputs from each iteration into the training set, continuously monitoring and enhancing the performance of the fine-tuned LLM to ensure its efficacy.
Knowledge graph construction
Given a collection of entities $E$ and relations $R$, a knowledge graph $\mathcal{K}=E\times R\times E$ is structured as a directed multi-relational graph. It comprises triples formatted as $(h,r,t)\in\mathcal{K}$, where $h$ and $t$ are entities within $E$. To structure the inference results into triples, we employ three labels that signify the material as the core label, with the core label serving as the head $h$, the names of other labels forming the relations $R$, and their values acting as tails $t$. Within the hierarchy of core labels, "Formula" is accorded the highest priority, followed by "Name", and then "Acronym". Furthermore, each head $h$ and tail $t$ node is linked to the DOI of the source article associated with the triplet. This setup allows us to ascertain the provenance of the relation between any two nodes by examining the intersection in their connected entities. Then we transfer the triples into MKG and store the MKG via graph database Neo4j, which also supports the subgraph matching function, where subgraph matching naturally suits the need to search for certain materials with user-input conditions. To facilitate access to the detailed information, we also make the dataset available in the RDF and CSV format for straightforward data handling.
Graph completion
In the analysis depicted in Figure 3, we enhance the traditional Jaccard similarity, hereafter referred to as $J$, to form a comprehensive similarity metric for evaluating potential material-to-application links. The Jaccard similarity is mathematically defined as:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
where $A$ and $B$ represent the attribute sets of two compared materials, such as their properties or applications. To refine our predication method, we extend this basic formulation by incorporating weighted factors that reflect the relevance and impact of each attribute, resulting in a modified similarity score for "Materials-Applications", $J_{\text{mat}}$. This score is computed as follows:
$J_{\text{mat}}(A,B)=\sum_{i=1}^{n}w_{i}\cdot\frac{|A_{i}\cap B_{i}|}{|A_{i}\cup B_{i}|}$
where $n$ is the number of attributes considered, $A_{i}$ and $B_{i}$ are the specific attributes of materials $A$ and $B$ respectively, and $w_{i}$ are the weights assigned to each attribute, signifying their importance in the context of material application. This approach mimics the investigative processes commonly employed by materials scientists when developing new materials. In determining the suitability of new materials for specific applications, researchers systematically analyze their chemical and physical properties, as well as their structural characteristics, comparing these attributes to those of materials well-established in the target field. This enhanced method not only accounts for the quantity of overlapping attributes but also emphasizes the significance of each attribute in relation to the material’s potential application. A higher $J_{\text{modified}}$ score indicates a stronger likelihood of applicability between the material in question and the target application. Thus, materials scoring high on this metric are considered promising candidates for the specified applications.
Figure 3: The process of network-based graph completion.
In addition to the Enhanced Jaccard Similarity, we employ the TransE model [27]. TransE treats entities and relations as vectors in the same embedding space. The core idea of TransE is to model relations by interpreting them as translations in the embedding space. For a given relation $r$ between head $h$ and tail $t$, the relationship is said to hold if:
$\|\mathbf{h}+\mathbf{r}-\mathbf{t}\|\approx 0$
[…]
Result
Figure 4: Schematic comparison of MKG and MatKG2.
To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG2 [15], highlighting the differences in their construction methodologies as illustrated in Figure 4. MatKG2 is built using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation. In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously. This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
To demonstrate the effectiveness of this pipeline, we evaluated the performance of each LLM on each task, as shown in the Table 1. Darwin significantly outperforms both the LLaMA 7b and LLaMA2 7b models in the F1 scores for NER and RE tasks, suggesting that Darwin yields more effective results in text-related tasks in materials science. However, there is no marked difference in the performance of these models on ER tasks, we can say LLM has a weaker ability to complete ER in our task. This may be attributed to the limited contextual memory capabilities of the LLMs and this is we need to apply additional ER process. The normailzed Darwin shows the performance of our ER and normalization process, the result indicate that it not only achieve the ER task successfully, but also contribute to the NER and RE task.
Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
Model Task Precision Recall F1 score
NER 0.6101 0.6216 0.6158
Llama 7b RE 0.5305 0.5405 0.5355
ER 0.3687 0.3757 0.3722
NER 0.7419 0.7667 0.7541
Llama2 7b RE 0.6452 0.6667 0.6557
ER 0.4484 0.4633 0.4557
NER 0.8013 0.7935 0.7974
Darwin RE 0.7036 0.6968 0.7002
ER 0.4593 0.4548 0.4571
NER 0.9520 0.9083 0.9296
Darwin (Normalized) RE 0.9039 0.8625 0.8827
ER 0.9127 0.8708 0.8913
To better understand the enhancements provided by each component in the ER/Normalization process, ablation studies were performed. The outcomes, detailed in Table 2, demonstrate that each technique positively affects the effectiveness of ER. The results highlight that utilizing the expert dictionary (ER-ED) is the most beneficial method, enhancing accuracy comprehensively across nearly all labels which are included in the dictionary. Additionally, the improvements seen with ER-NF/A are noteworthy; this step is based on the ChemDataExactor, effectively removes the majority of incorrect material identifications in core labels and offers even more substantial benefits to NER than to ER-ED. Furthermore, from the results of the ablation experiment, ER-N/F appears that its contribution is not as significant as the other two components, but considering the confusion between the material Formula and Name may have an impact on subsequent MKG applications. ER-N/F is also an indispensable part of the ER process. The normalized data is transformed into triples to construct the MKG, which comprises 162,605 nodes and 731,772 edges, as illustrated in the schematic diagram in Figure 5.
Table 2: Result of the ablation experiment in normalization.
  Method NER F1 $\nabla$ RE F1 $\nabla$ ER F1 $\nabla$
Darwin (Normalized) 92.96 - 88.27 - 89.13 -
ER-N/F 92.96 - 83.29 -4.98 89.13 -
ER-NF/A 84.01 -8.85 83.07 -5.20 84.54 -4.59
ER-ED 88.59 -4.37 80.20 -8.07 50.00 -38.83
Figure 5: (a) Global schematic diagram of MKG; (b) Local schematic diagram of MKG.
Finally, to demonstrate the reliability and accuracy of our similarity algorithm, we divided MKG into two knowledge graphs based on publication year of paper. KG with earlier years applied similarity calculation algorithms to predict potential links, while KG with later years are used to verify how many sets of "Material Applications" in these predictions were reported in the following n years. The result shows in Figure 6. Specifically, in Figure 6 (a), the grey line uses only abstracts published before the year to make predication and the percentage of reported predication in next few year is displayed. The predictive capacity of the MKG is substantial, therefore, for each prediction, we only use the top 200 "material-application" pairs for validation. Looking at the results, as time progresses, the predicted materials are gradually verified. Using data from 2014, within nine years, 48.5% of the "material-application" predictions have been validated. This outcome demonstrates the effectiveness and reliability of the MKG, and it also lays the foundation for proposing new materials in some fields. The red, blue and green line are average percentage of reported link based on network similarity, Jaccard similarity directly between "Formula" and "Application", and TransE for link predication, respectively. Network-based similarity has the highest accurate predication which indicates the effective of simulating the thinking of real material development using algorithms. The Figure 6 (b) is an example, visualizing the KG made from data before 2018, and the predicted "Material Application" reported within 5 years using network based algorithms.
Figure 6: Validation of the graph completion. (a) Percentage of reported predication after years. (b) Example of predicated material - application using data before 2018.
Conclusion and future work
In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction aimed at efficiently extracting triples from unstructured scientific texts. This methodology distinguishes itself by enabling the fine-tuning of Large Language Models (LLMs) with minimally annotated datasets. Subsequently, the fine-tuned LLMs are utilized to extract structured information from extensive corpora of unstructured text. This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted.
Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers. Our analysis confirms the effectiveness and credibility of the MKG. Furthermore, we have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding. This analysis has enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions. First, by extending the depth of structured information extraction to cover entire research papers, we anticipate the development of a richer and more detailed KG. This extension involves not merely enlarging the dataset but also enhancing our extraction techniques to accurately capture subtle distinctions within complex scientific literature. Second, improving the precision of entity labeling in our pipeline enables more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties. This enhancement significantly increases the granularity and practical utility of the KG. Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science. Lastly, the integration of our Material Knowledge Graph with existing KGs, such as MatKG, paves the way for the creation of a more interconnected and expansive dataset. This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.
Appendix
A. Validation for MKG
To validation of MKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science. After splitting these triples apart, annotator can obtain 1000 entities and 500 relations. The report from annotator shows in Table 3. The labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation. This is due to the rigorous normalization of entities under these categories, using our expert dictionary. However, strict normalization can also lead to certain entity loss, which is difficult to avoid. In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation. But the results are still satisfactory. "Name" and "Acronym" has also achieved a commendable 100% entity accuracy. However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations. A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor. Fortunately, the impact of this kind of error is not significant and will not bring a significant influence, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
Table 3: Human evaluation metric on randomly selected triples.
Label Number in total Entity disagree % disagree Relation disagree % disagree
Formula 257 0 0 N/A N/A
Name 227 0 0 12 5.3
Acronym 26 0 0 3 11.5
Descriptor 127 7 5.5 12 9.4
Property 147 12 8.2 9 6.1
Application 73 0 0 0 0
Structure/Phase 36 0 0 0 0
Synthesis 50 0 0 0 0
Characterization 57 0 0 0 0
B. Predication result
As Figure 7 illustrates, we display the materials ranked in the top five for 2023 across various domains, alongside their rankings in previous years based on available data up to 2023. It is important to note that these materials-application pairs were not ranked within the top five before 2017, which is why they do not appear in the figure for those earlier years. Stars mark the years these materials were featured in literature for their respective applications. The figure emphasizes the significance and emerging utility of these materials by highlighting the top five in applications such as catalysts, batteries, fuel/solar cells, and other fields from 2017 to 2023, indicating their growing importance and potential to revolutionize technology in these areas. This predication serves as an inspiration to materials researchers, accelerating progress in material development. For example, consider the material NiFeN HC/NF; it not only excels in the Oxygen Evolution Reaction (OER) but has also demonstrated its capability as a catalyst for the Hydrogen Evolution Reaction (HER). This means that NiFeN HC/NF is a promising bifunctional electrocatalyst for water electrolysis, aligning with our understanding of its potential in dual-function applications.
C. Paramaters for LLM fine tuning
epochs: 10 train batch size (per device): 1 eval batch size (per device): 1 gradient accumulation steps: 2 evaluation strategy: "no" save strategy: "steps" save steps: 500 save total limit: 1 learning rate: 2e-5 weight decay: 0. warmup ratio: 0.03 lr scheduler type: "cosine" logging steps: 1 tf32: False fsdp: "full shard auto wrap" fsdp transformer layer cls to wrap: "LlamaDecoderLayer" token max length: 1024
Figure 7: High rank predicated materials in (a) Catalyst, (b) Fuel/Solar cell, (c) Battery and (d) Other fields, the stars represent the material has been reported in this application.
D. Experiments compute resources
The compute resources were primarily consumed in LLM fine-tuning, involving the use of 8 GPUs for 9.02 hours with a 36% utilization rate, consuming 25.98 service units and using 14245 MiB of storage from a quota of 81920 MiB.